OpenStudy (chazz16):

NEED HELP MEDAL WILL BE AWARDED (y-4) • (y+8)

OpenStudy (anonymous):

whats your question?

OpenStudy (anonymous):

What do you need help with?

OpenStudy (chazz16):

(y-4) • (y+8)

OpenStudy (texaschic101):

are you familiar with the FOIL method

OpenStudy (anonymous):

do you want the answer or how to do it?

OpenStudy (chazz16):

Not At All I Don't Get This At All.

OpenStudy (anonymous):

I know the answer hold on let me type it!

OpenStudy (chazz16):

I Don't Care How Its Given To Me I Just Need To Know The Answer.

OpenStudy (anonymous):

Reorder the terms: 4(-4 + y) = 8 (-4 * 4 + y * 4) = 8 (-16 + 4y) = 8 Solving -16 + 4y = 8 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '16' to each side of the equation. -16 + 16 + 4y = 8 + 16 Combine like terms: -16 + 16 = 0 0 + 4y = 8 + 16 4y = 8 + 16 Combine like terms: 8 + 16 = 24 4y = 24 Divide each side by '4'. y = 6 Simplifying y = 6

OpenStudy (anonymous):

=y^2+4y-32

OpenStudy (texaschic101):

FOIL -- first, outside,inside,last (y - 4)(y + 8) F -- first multiply the first terms in each set....y * y = y^2 O -- multiply outside terms in each set...y * 8 = 8y I -- multiply inside terms in each set...-4 * y = -4y L -- multiply last terms in each set...-4 * 8 = -32 now put it all together... y^2 + 8y - 4y - 32 -- combine like terms y^2 + 4y - 32

OpenStudy (chazz16):

Great Thanks I Have More Questions Medals Will Be Awarded

OpenStudy (anonymous):

alright

OpenStudy (chazz16):

Next Question Need Help Medal Will Be Awarded Hears The Question Link http://openstudy.com/study#/updates/56279ea5e4b09cb8ec244227